OpenStudy (anonymous):

find the sixtieth term of the arithmetic sequence if a1=418 and d = 12

OpenStudy (anonymous):

tn=a+(n-1)d

OpenStudy (anonymous):

@surjithayer thank you

OpenStudy (imstuck):

\[a _{60}=418+(60-1)12--->a _{60}=418+[59\times12]--->a _{60}=418+708\] 418 + 708= 1126

OpenStudy (anonymous):

yw